# Rui Mao1, Chao Huang1, Jing Bi1, Zeliang Zhang1, Pooyan Fazli2, Chenliang Xu1†

### v4

@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) [1, 44, 28, 9, 48, 16, 3] have greatly enhanced capabilities in understanding multimodality. However, while current benchmarks [11, 34, 26, 13] for evaluating MLLMs assess general image or video comprehension, they lack a detailed focus on video composition, the nuanced interpretation of how visual elements combine and interact within compiled videos. Compiled videos refer to those created by editing and integrating multiple clips, scenes, or sequences, either from various sources or from different segments of a single recording, e.g. films, TV series, documentaries, animations, vlogs, etc. These videos are carefully constructed to create a seamless flow and include richer compositions, requiring shot-by-shot analysis to interpret.
@@
Through investigating existing benchmarks, we identified their limitations in evaluating MLLM in video composition understanding. As shown in Table 1, the benchmarks in the first group [11, 32, 17] primarily focus on static images and overlook the dynamic aspects of visual content. Among these, Winoground [49] and MMComposition [17] attempts to assess the compositionality of MLLMs, though it is limited to image-based evaluations. The second group consists of traditional video benchmarks [55, 22, 59, 54], which are less effective at addressing the specific limitations of modern MLLMs. While TVQA [22] includes a compositional video QA component, its compositionality is relatively coarse-grained, limited to basic question types like “who,” “when,” “where,” “how,” and “what.” The third group highlights recent benchmarks [6, 26, 41, 45, 33, 51, 13] developed to assess MLLMs’ video comprehension capabilities. Although these benchmarks incorporate tasks that touch on compositional understanding, their evaluations of compositionality remain limited. Additionally, most videos in these benchmarks are natural-shot rather than compiled, posing a challenge for models trained on natural footage to effectively interpret the increasingly prevalent edited and compiled videos seen on modern online video platforms.
@@
We evaluate 33 state-of-the-art MLLMs on VidComposition, including 27 open-source and 6 proprietary models, revealing a substantial performance gap between MLLMs and human-level comprehension in video composition understanding. As shown in Figure 1, although top models [24, 8, 50, 9, 1, 44, 29] perform well on basic perception tasks (e.g. action perception), they fall short in comprehending complex video compositions, particularly in cinematography. This performance disparity underscores current models’ limitations in capturing intricate, multi-layered video structures. Additional experiments further analyze factors influencing MLLM performance, such as the number of frames provided as input, the resolution of visual encoders, the size of language decoders, and the data volume for fine-tuning, yielding insights for future advancements in model design. Overall, our benchmark offers valuable insights for enhancing MLLMs and also suggests applications in video generation where MLLMs could assist in automatically evaluating the compositional quality of generated videos.
@@
Winoground [49] I 400 8 - M
MME [11] I 1.1k 14 - M
MMBench [34] I 1.7k 20 - A+M
MMComposition [17] I 4.3k 13 - M
MSVD-QA [55] V 504 5 A
MSRVTT-QA [55] V 2.9k 5 A
TGIF-QA [19] V 9.6k 4 A
TVQA [22] V 2.2k 8 A+M
ActivityNet-QA [59] V 5.8k 4 M
NExT-QA [54] V 1k 8 A
AutoEval-Video [6] V 327 9 A+M
Video-Bench [41] V 5.9k 10 A+M
LVBench [51] V 500 6 M
MVBench [26] V 3.6k 20 A+M
MovieChat-1k [45] V 100 8 M
TempCompass [33] V 410 4 M
Video-MME [13] V 900 12 M
@@ 2.3 Evaluation Metrics
LLaVA-OneVision-72B [24] 57.1 60.5 66.3 63.3 90.0 90.0 74.6 84.7 72.4 76.9 12.0 90.3 89.5 34.5 74.1 63.31
InternVL2-40B [7] 46.6 60.0 58.9 51.0 90.0 83.3 67.6 79.6 51.9 80.0 47.4 64.5 68.4 44.8 85.2 60.73
InternVL2-76B [7] 46.6 63.9 45.5 51.0 86.7 86.7 76.1 81.3 49.0 80.0 44.5 51.6 76.3 24.1 75.9 58.73
Qwen2-VL-72B [50] 37.9 56.6 57.9 34.7 76.7 76.7 63.4 76.2 66.5 73.8 53.6 74.2 65.8 3.4 55.6 58.68
Video-LLaMA2-72B [9] 47.5 56.1 59.4 63.3 76.7 80.0 71.8 68.5 63.2 73.8 36.8 74.2 60.5 34.5 72.2 58.62
InternVL2-8B [7] 55.3 56.6 59.4 44.9 80.0 83.3 59.2 67.2 40.6 78.5 32.5 64.5 52.6 31.0 72.2 54.63
GPT-4o [1] 40.2 37.1 59.4 51.0 73.3 90.0 40.8 90.6 41.4 72.3 27.8 51.6 81.6 34.5 77.8 52.93
Gemini-1.5-Flash [44] 43.4 32.7 52.0 55.1 70.0 48.4 62.0 78.7 47.3 83.1 26.3 71.0 78.9 44.8 70.4 52.40
VILA-1.5-40B [29] 32.0 56.6 54.5 42.9 83.3 86.7 57.7 67.7 44.4 75.4 22.5 74.2 65.8 48.3 77.8 51.23
GPT-4o mini [1] 33.8 49.8 50.5 49.0 80.0 90.0 31.0 79.6 41.4 66.2 26.8 61.3 76.3 20.7 79.6 50.23
Gemini-1.5-Pro [44] 33.8 51.7 51.5 51.0 73.3 80.0 70.4 47.7 36.4 73.8 37.3 71.0 84.2 58.6 75.9 49.36
Qwen2-VL-7B [50] 20.1 46.8 37.1 38.8 70.0 76.7 54.9 73.2 49.4 72.3 52.2 61.3 42.1 17.2 70.4 49.30
Oryx-7B [35] 34.7 54.1 57.4 57.1 80.0 73.3 66.2 48.5 34.7 73.8 41.6 61.3 39.5 20.7 66.7 48.77
Gemini-1.5-Flash-8B [44] 43.4 45.9 56.9 36.7 70.0 76.7 35.2 69.8 36.0 73.8 26.8 48.4 71.1 24.1 64.8 48.59
Video-LLaMA2.1 [9] 44.3 35.6 39.6 51.0 76.7 83.3 50.7 60.9 45.2 75.4 35.4 58.1 36.8 20.7 81.5 47.77
VideoChat2 [26] 24.2 58.0 42.1 44.9 66.7 83.3 60.6 62.6 27.6 73.8 55.0 35.5 50.0 10.3 59.3 47.36
InternVL2-26B [7] 33.3 47.8 39.6 55.1 76.7 83.3 56.3 68.9 33.9 76.9 25.4 45.2 34.2 41.4 75.9 46.42
LongVA [62] 24.7 41.0 48.0 40.8 70.0 73.3 42.3 51.9 32.2 72.3 42.6 48.4 52.6 34.5 70.4 43.73
MiniCPM-V2.6 [56] 28.3 43.4 43.6 53.1 73.3 80.0 50.7 59.1 23.0 75.4 22.0 71.0 57.9 20.7 72.2 42.49
InternVL2-4B [7] 27.4 42.9 26.2 32.7 66.7 73.3 49.3 60.4 28.0 78.5 41.6 35.5 44.7 10.3 72.2 41.68
Video-LLaMA2.1-AV [9] 27.4 45.9 38.6 55.1 73.3 76.7 47.9 46.4 30.1 81.5 25.8 45.2 34.2 34.5 83.3 41.50
VILA-1.5-8B [29] 31.5 40.0 37.6 51.0 63.3 66.7 40.8 40.9 26.8 70.8 37.8 41.9 60.5 44.8 59.3 40.21
GPT-4-turbo [1] 23.7 37.1 35.1 46.9 63.3 80.0 25.4 54.9 36.4 50.8 29.7 64.5 39.5 44.8 70.4 39.85
LongLLaVA [53] 28.3 37.1 27.2 24.5 60.0 56.7 54.9 48.1 32.6 61.5 38.3 41.9 26.3 24.1 66.7 38.45
Kangaroo [31] 29.2 42.0 24.3 30.6 56.7 66.7 57.7 31.5 26.8 67.7 47.8 61.3 21.1 6.9 55.6 37.10
InternVL2-2B [7] 23.7 24.4 24.8 36.7 76.7 63.3 53.5 48.9 21.8 80.0 40.2 29.0 47.4 6.9 83.3 36.75
LongVILA [29] 25.1 35.6 40.6 40.8 80.0 60.0 38.0 32.8 25.1 76.9 20.6 64.5 50.0 37.9 79.6 36.46
AuroraCap [5] 34.7 42.0 40.6 46.9 53.3 56.7 35.2 35.3 22.2 63.1 22.5 32.3 50.0 31.0 59.3 36.28
Qwen2-VL-2B [50] 21.0 29.3 25.2 30.6 63.3 70.0 42.3 50.6 23.8 67.7 37.3 74.2 34.2 24.1 63.0 36.16
Video-LLaMA2-7B [9] 25.1 29.3 23.3 30.6 70.0 66.7 40.8 31.5 26.4 72.3 40.2 29.0 44.7 27.6 66.7 34.35
VILA-1.5-3B [29] 20.1 32.7 38.1 51.0 53.3 46.7 31.0 26.4 10.5 72.3 35.4 32.3 36.8 10.3 83.3 31.95
Video-LLaVA [28] 26.5 25.9 38.1 32.7 53.3 40.0 25.4 26.8 23.0 55.4 30.1 38.7 21.1 51.7 51.9 31.07
Chat-UniVi [20] 25.1 31.7 30.2 30.6 53.3 30.0 22.5 26.4 24.3 29.2 21.1 32.3 34.2 44.8 40.7 28.02
InternVL2-1B [7] 24.7 24.9 22.8 22.4 46.7 33.3 22.5 26.4 26.8 30.8 30.6 22.6 23.7 34.5 29.6 26.61
@@ 3 Main Results
Overall Performance. As shown in Table 2, the overall performance on the VidComposition benchmark reveals that understanding intricate video compositions remains challenging for MLLMs. While humans achieve exceptionally high scores (86.26), the leading models, such as LLaVA-OneVision-72B [24] (63.31), InternVL2-40B [7] (60.73), InternVL2-76B [7] (58.7) and Qwen2-VL-72B [50] (58.78), demonstrate only moderate success, underscoring the complexity of the tasks and the current limitations in video composition understanding. This gap between human and model performance highlights the benchmark’s rigor and the need for advancements in fine-grained video-based compositional learning. Open-source models with advanced vision components, particularly InternVL2 variants, outperform API-based models like GPT-4o [1] (52.93) and Genmini-1.5-Flash [44] (52.40). The mean overall accuracy of these MLLMs is 43.44. For reference, the random-choice baseline has a score of 25.33. While the models exceed it, they still face considerable obstacles in approaching human-level video composition understanding.
@@
Chat-UniVi [20]; Video-LLaVA [28]; VideoChat2 [26] 8 7B 224 31.68 42.22 31.02 40.11 42.98 34.94
Chat-UniVi-v1.5 [20]; LongVA [62]; Video-LLaMA2 [9] 336 33.6+1.92 48.89+6.67 32.42+1.4 44.26+4.15 50.96+7.98 38.04+3.1
Video-LLaMA2.1 [9]; Video-LLaMA2.1-AV [9] 384 36.9+3.3 55.28+6.39 46.0+13.58 43.11-1.15 53.72+2.76 43.7+5.66
Chat-UniVi [20]; Video-LLaVA [28]; VideoChat2 [26] 16 7B 224 29.66 39.81 31.86 28.52 34.71 31.52
Chat-UniVi-v1.5 [20]; LongVA [62]; Video-LLaMA2 [9] 336 32.75+3.09 49.07+9.26 32.21+0.35 44.92+16.4 49.04+14.33 37.67+6.15
Video-LLaMA2.1 [9]; Video-LLaMA2.1-AV [9] 384 37.06+4.31 57.22+8.15 45.68+13.47 43.44-1.48 54.13+5.09 43.96+6.29
LongVA [62]; Video-LLaMA2 [9] 32 7B 336 31.39 46.67 35.26 44.26 53.72 37.98
Video-LLaMA2.1 [9]; Video-LLaMA2.1-AV [9] 384 38.5+7.11 59.72+13.05 45.47+10.21 42.95-1.31 54.55+0.83 44.64+6.66
Strengths & Weaknesses Analysis. From Table 2, we see that MLLMs generally perform better in CU tasks, particularly A-P and CMP-P. For example, top models like LLaVA-OneVision-72B [24], GPT-4o [1] and GPT-4o mini [1] achieve high scores in CMP-P; LLaVA-OneVision-72B [24] and InternVL2-40B [7] get 90.0 on A-P. This indicates that state-of-the-art MLLMs can effectively recognize and interpret actions and visual details of characters in a scene. Models also show strong performance on SP tasks such as B-P and L-P, with models like Gemini-1.5-Flash [44] achieves 83.1 on B-P and LLaVA-OneVision-72B [24] gets 90.3 on L-P. Additionally, these models achieve competitive scores on some MA tasks, such as AS-P and SE-P, with top scores reaching 89.5 and 85.2, respectively. This strong performance may be attributed to the fact that these tasks rely on some expert knowledge about video-making techniques, which MLLMs can acquire from massive corpora. Conversely, the models encounter significant difficulties in more complex compositional tasks, especially CA. For example, CamM-P and SS-P yield only modest scores, with top models reaching 57.1 and 63.9, respectively, reflecting a significant gap in understanding cinematic techniques. NU tasks, such as S-M and P-O, also present challenges, with model performance substantially trailing behind human benchmarks because there is often a gap between scripts and actual video presentation. Unlike humans, who can intuitively bridge this gap, MLLMs are fine-tuned on closely matched vision-text pairs, limiting their ability to interpret subtle or implied connections in narrative tasks. Additionally, counting tasks, such as Cha-C, S-C, and Cut-C, remain particularly problematic for most models, further underscoring the limitations in visual counting abilities and understanding scene transitions across multiple frames across all models.

### v5

@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) [achiam2023gpt, reid2024gemini, lin2023videollava, cheng2024videollama2, tang2024avicuna, hua2024v2xum, bi2024eagle] have greatly enhanced capabilities in understanding multimodality. However, while current benchmarks [fu2023mme, liu2025mmbench, li2024mvbench, fu2024video-mme] for evaluating MLLMs assess general image or video comprehension, they lack a detailed focus on video composition, the nuanced interpretation of how visual elements combine and interact within compiled videos. Compiled videos refer to those created by editing and integrating multiple clips, scenes, or sequences, either from various sources or from different segments of a single recording, e.g. films, TV series, documentaries, animations, vlogs, etc. These videos are carefully constructed to create a seamless flow and include richer compositions, requiring shot-by-shot analysis to interpret.
@@
Through investigating existing benchmarks, we identified their limitations in evaluating MLLM in video composition understanding. As shown in Table 1, the benchmarks in the first group [fu2023mme, liu2023mmbench, hua2024mmcomposition] primarily focus on static images and overlook the dynamic aspects of visual content. Among these, Winoground [thrush2022winoground] and MMComposition [hua2024mmcomposition] attempts to assess the compositionality of MLLMs, though it is limited to image-based evaluations. The second group consists of traditional video benchmarks [xu2017video, lei2018tvqa, yu2019activitynet, NExT-QA], which are less effective at addressing the specific limitations of modern MLLMs. While TVQA [lei2018tvqa] includes a compositional video QA component, its compositionality is relatively coarse-grained, limited to basic question types like “who,” “when,” “where,” “how,” and “what.” The third group highlights recent benchmarks [chen2023autoevalvideo, li2024mvbench, ning2023video-bench, song2024moviechat, liu2024tempcompass, wang2024lvbench, fu2024video-mme] developed to assess MLLMs’ video comprehension capabilities. Although these benchmarks incorporate tasks that touch on compositional understanding, their evaluations of compositionality remain limited. Additionally, most videos in these benchmarks are natural-shot rather than compiled, posing a challenge for models trained on natural footage to effectively interpret the increasingly prevalent edited and compiled videos seen on modern online video platforms.
@@
We evaluate 33 state-of-the-art MLLMs on VidComposition, including 27 open-source and 6 proprietary models, revealing a substantial performance gap between MLLMs and human-level comprehension in video composition understanding. As shown in Figure 1, although top models [li2024llavaonevisioneasyvisualtask, chen2024internvl, Qwen2-VL, cheng2024videollama2, achiam2023gpt, reid2024gemini, lin2024vila] perform well on basic perception tasks (e.g. action perception), they fall short in comprehending complex video compositions, particularly in cinematography. This performance disparity underscores current models’ limitations in capturing intricate, multi-layered video structures. Additional experiments further analyze factors influencing MLLM performance, such as the number of frames provided as input, the resolution of visual encoders, the size of language decoders, and the data volume for fine-tuning, yielding insights for future advancements in model design. Overall, our benchmark offers valuable insights for enhancing MLLMs and also suggests applications in video generation where MLLMs could assist in automatically evaluating the compositional quality of generated videos.
@@
Winoground [thrush2022winoground] I 400 8 - M
MME [fu2023mme] I 1.1k 14 - M
MMBench [liu2025mmbench] I 1.7k 20 - A+M
MMComposition [hua2024mmcomposition] I 4.3k 13 - M
MSVD-QA [xu2017video] V 504 5 A
MSRVTT-QA [xu2017video] V 2.9k 5 A
TGIF-QA [jang2017tgif] V 9.6k 4 A
TVQA [lei2018tvqa] V 2.2k 8 A+M
ActivityNet-QA [yu2019activitynet] V 5.8k 4 M
NExT-QA [NExT-QA] V 1k 8 A
AutoEval-Video [chen2023autoevalvideo] V 327 9 A+M
Video-Bench [ning2023video-bench] V 5.9k 10 A+M
LVBench [wang2024lvbench] V 500 6 M
MVBench [li2024mvbench] V 3.6k 20 A+M
MovieChat-1k [song2024moviechat] V 100 8 M
TempCompass [liu2024tempcompass] V 410 4 M
Video-MME [fu2024video-mme] V 900 12 M
@@ 2.3 Evaluation Metrics
LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask] 57.1 60.5 66.3 63.3 90.0 90.0 74.6 84.7 72.4 76.9 12.0 90.3 89.5 34.5 74.1 63.31
InternVL2-40B [chen2024far] 46.6 60.0 58.9 51.0 90.0 83.3 67.6 79.6 51.9 80.0 47.4 64.5 68.4 44.8 85.2 60.73
InternVL2-76B [chen2024far] 46.6 63.9 45.5 51.0 86.7 86.7 76.1 81.3 49.0 80.0 44.5 51.6 76.3 24.1 75.9 58.73
Qwen2-VL-72B [Qwen2-VL] 37.9 56.6 57.9 34.7 76.7 76.7 63.4 76.2 66.5 73.8 53.6 74.2 65.8 3.4 55.6 58.68
Video-LLaMA2-72B [cheng2024videollama2] 47.5 56.1 59.4 63.3 76.7 80.0 71.8 68.5 63.2 73.8 36.8 74.2 60.5 34.5 72.2 58.62
InternVL2-8B [chen2024far] 55.3 56.6 59.4 44.9 80.0 83.3 59.2 67.2 40.6 78.5 32.5 64.5 52.6 31.0 72.2 54.63
GPT-4o [achiam2023gpt] 40.2 37.1 59.4 51.0 73.3 90.0 40.8 90.6 41.4 72.3 27.8 51.6 81.6 34.5 77.8 52.93
Gemini-1.5-Flash [reid2024gemini] 43.4 32.7 52.0 55.1 70.0 48.4 62.0 78.7 47.3 83.1 26.3 71.0 78.9 44.8 70.4 52.40
VILA-1.5-40B [lin2024vila] 32.0 56.6 54.5 42.9 83.3 86.7 57.7 67.7 44.4 75.4 22.5 74.2 65.8 48.3 77.8 51.23
GPT-4o mini [achiam2023gpt] 33.8 49.8 50.5 49.0 80.0 90.0 31.0 79.6 41.4 66.2 26.8 61.3 76.3 20.7 79.6 50.23
Gemini-1.5-Pro [reid2024gemini] 33.8 51.7 51.5 51.0 73.3 80.0 70.4 47.7 36.4 73.8 37.3 71.0 84.2 58.6 75.9 49.36
Qwen2-VL-7B [Qwen2-VL] 20.1 46.8 37.1 38.8 70.0 76.7 54.9 73.2 49.4 72.3 52.2 61.3 42.1 17.2 70.4 49.30
Oryx-7B [liu2024oryxmllmondemandspatialtemporal] 34.7 54.1 57.4 57.1 80.0 73.3 66.2 48.5 34.7 73.8 41.6 61.3 39.5 20.7 66.7 48.77
Gemini-1.5-Flash-8B [reid2024gemini] 43.4 45.9 56.9 36.7 70.0 76.7 35.2 69.8 36.0 73.8 26.8 48.4 71.1 24.1 64.8 48.59
Video-LLaMA2.1 [cheng2024videollama2] 44.3 35.6 39.6 51.0 76.7 83.3 50.7 60.9 45.2 75.4 35.4 58.1 36.8 20.7 81.5 47.77
VideoChat2 [li2024mvbench] 24.2 58.0 42.1 44.9 66.7 83.3 60.6 62.6 27.6 73.8 55.0 35.5 50.0 10.3 59.3 47.36
InternVL2-26B [chen2024far] 33.3 47.8 39.6 55.1 76.7 83.3 56.3 68.9 33.9 76.9 25.4 45.2 34.2 41.4 75.9 46.42
LongVA [zhang2024longva] 24.7 41.0 48.0 40.8 70.0 73.3 42.3 51.9 32.2 72.3 42.6 48.4 52.6 34.5 70.4 43.73
MiniCPM-V2.6 [yao2024minicpm] 28.3 43.4 43.6 53.1 73.3 80.0 50.7 59.1 23.0 75.4 22.0 71.0 57.9 20.7 72.2 42.49
InternVL2-4B [chen2024far] 27.4 42.9 26.2 32.7 66.7 73.3 49.3 60.4 28.0 78.5 41.6 35.5 44.7 10.3 72.2 41.68
Video-LLaMA2.1-AV [cheng2024videollama2] 27.4 45.9 38.6 55.1 73.3 76.7 47.9 46.4 30.1 81.5 25.8 45.2 34.2 34.5 83.3 41.50
VILA-1.5-8B [lin2024vila] 31.5 40.0 37.6 51.0 63.3 66.7 40.8 40.9 26.8 70.8 37.8 41.9 60.5 44.8 59.3 40.21
GPT-4-turbo [achiam2023gpt] 23.7 37.1 35.1 46.9 63.3 80.0 25.4 54.9 36.4 50.8 29.7 64.5 39.5 44.8 70.4 39.85
LongLLaVA [wang2024longllava] 28.3 37.1 27.2 24.5 60.0 56.7 54.9 48.1 32.6 61.5 38.3 41.9 26.3 24.1 66.7 38.45
Kangaroo [liu2024kangaroo] 29.2 42.0 24.3 30.6 56.7 66.7 57.7 31.5 26.8 67.7 47.8 61.3 21.1 6.9 55.6 37.10
InternVL2-2B [chen2024far] 23.7 24.4 24.8 36.7 76.7 63.3 53.5 48.9 21.8 80.0 40.2 29.0 47.4 6.9 83.3 36.75
LongVILA [lin2024vila] 25.1 35.6 40.6 40.8 80.0 60.0 38.0 32.8 25.1 76.9 20.6 64.5 50.0 37.9 79.6 36.46
AuroraCap [chai2024auroracap] 34.7 42.0 40.6 46.9 53.3 56.7 35.2 35.3 22.2 63.1 22.5 32.3 50.0 31.0 59.3 36.28
Qwen2-VL-2B [Qwen2-VL] 21.0 29.3 25.2 30.6 63.3 70.0 42.3 50.6 23.8 67.7 37.3 74.2 34.2 24.1 63.0 36.16
Video-LLaMA2-7B [cheng2024videollama2] 25.1 29.3 23.3 30.6 70.0 66.7 40.8 31.5 26.4 72.3 40.2 29.0 44.7 27.6 66.7 34.35
VILA-1.5-3B [lin2024vila] 20.1 32.7 38.1 51.0 53.3 46.7 31.0 26.4 10.5 72.3 35.4 32.3 36.8 10.3 83.3 31.95
Video-LLaVA [lin2023videollava] 26.5 25.9 38.1 32.7 53.3 40.0 25.4 26.8 23.0 55.4 30.1 38.7 21.1 51.7 51.9 31.07
Chat-UniVi [jin2024chat-univi] 25.1 31.7 30.2 30.6 53.3 30.0 22.5 26.4 24.3 29.2 21.1 32.3 34.2 44.8 40.7 28.02
InternVL2-1B [chen2024far] 24.7 24.9 22.8 22.4 46.7 33.3 22.5 26.4 26.8 30.8 30.6 22.6 23.7 34.5 29.6 26.61
@@ 3 Main Results
Overall Performance. As shown in Table 2, the overall performance on the VidComposition benchmark reveals that understanding intricate video compositions remains challenging for MLLMs. While humans achieve exceptionally high scores (86.26), the leading models, such as LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask] (63.31), InternVL2-40B [chen2024far] (60.73), InternVL2-76B [chen2024far] (58.7) and Qwen2-VL-72B [Qwen2-VL] (58.78), demonstrate only moderate success, underscoring the complexity of the tasks and the current limitations in video composition understanding. This gap between human and model performance highlights the benchmark’s rigor and the need for advancements in fine-grained video-based compositional learning. Open-source models with advanced vision components, particularly InternVL2 variants, outperform API-based models like GPT-4o [achiam2023gpt] (52.93) and Genmini-1.5-Flash [reid2024gemini] (52.40). The mean overall accuracy of these MLLMs is 43.44. For reference, the random-choice baseline has a score of 25.33. While the models exceed it, they still face considerable obstacles in approaching human-level video composition understanding.
@@
Chat-UniVi [jin2024chat-univi]; Video-LLaVA [lin2023videollava]; VideoChat2 [li2024mvbench] 8 7B 224 31.68 42.22 31.02 40.11 42.98 34.94
Chat-UniVi-v1.5 [jin2024chat-univi]; LongVA [zhang2024longva]; Video-LLaMA2 [cheng2024videollama2] 336 33.6+1.92 48.89+6.67 32.42+1.4 44.26+4.15 50.96+7.98 38.04+3.1
Video-LLaMA2.1 [cheng2024videollama2]; Video-LLaMA2.1-AV [cheng2024videollama2] 384 36.9+3.3 55.28+6.39 46.0+13.58 43.11-1.15 53.72+2.76 43.7+5.66
Chat-UniVi [jin2024chat-univi]; Video-LLaVA [lin2023videollava]; VideoChat2 [li2024mvbench] 16 7B 224 29.66 39.81 31.86 28.52 34.71 31.52
Chat-UniVi-v1.5 [jin2024chat-univi]; LongVA [zhang2024longva]; Video-LLaMA2 [cheng2024videollama2] 336 32.75+3.09 49.07+9.26 32.21+0.35 44.92+16.4 49.04+14.33 37.67+6.15
Video-LLaMA2.1 [cheng2024videollama2]; Video-LLaMA2.1-AV [cheng2024videollama2] 384 37.06+4.31 57.22+8.15 45.68+13.47 43.44-1.48 54.13+5.09 43.96+6.29
LongVA [zhang2024longva]; Video-LLaMA2 [cheng2024videollama2] 32 7B 336 31.39 46.67 35.26 44.26 53.72 37.98
Video-LLaMA2.1 [cheng2024videollama2]; Video-LLaMA2.1-AV [cheng2024videollama2] 384 38.5+7.11 59.72+13.05 45.47+10.21 42.95-1.31 54.55+0.83 44.64+6.66
Strengths & Weaknesses Analysis. From Table 2, we see that MLLMs generally perform better in CU tasks, particularly A-P and CMP-P. For example, top models like LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask], GPT-4o [achiam2023gpt] and GPT-4o mini [achiam2023gpt] achieve high scores in CMP-P; LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask] and InternVL2-40B [chen2024far] get 90.0 on A-P. This indicates that state-of-the-art MLLMs can effectively recognize and interpret actions and visual details of characters in a scene. Models also show strong performance on SP tasks such as B-P and L-P, with models like Gemini-1.5-Flash [reid2024gemini] achieves 83.1 on B-P and LLaVA-OneVision-72B [li2024llavaonevisioneasyvisualtask] gets 90.3 on L-P. Additionally, these models achieve competitive scores on some MA tasks, such as AS-P and SE-P, with top scores reaching 89.5 and 85.2, respectively. This strong performance may be attributed to the fact that these tasks rely on some expert knowledge about video-making techniques, which MLLMs can acquire from massive corpora. Conversely, the models encounter significant difficulties in more complex compositional tasks, especially CA. For example, CamM-P and SS-P yield only modest scores, with top models reaching 57.1 and 63.9, respectively, reflecting a significant gap in understanding cinematic techniques. NU tasks, such as S-M and P-O, also present challenges, with model performance substantially trailing behind human benchmarks because there is often a gap between scripts and actual video presentation. Unlike humans, who can intuitively bridge this gap, MLLMs are fine-tuned on closely matched vision-text pairs, limiting their ability to interpret subtle or implied connections in narrative tasks. Additionally, counting tasks, such as Cha-C, S-C, and Cut-C, remain particularly problematic for most models, further underscoring the limitations in visual counting abilities and understanding scene transitions across multiple frames across all models.